A 20.0-kg horizontal plank 4.00 m long rests on two supports, one at the left end and a second 1 .00 m from the right end. What is the magnitude of the force exerted on the plank by the support near the right end?
(d) 131 N.
(c) 112 N
(b) 45.2 N
(a) 32.0
1
Expert's answer
2020-12-03T10:59:46-0500

"L=4m"

mass of plank, "m=20kg"

take counterclockwise torques are positive

Apply torque equation about one end

"F_{1}(0) - mg(L\/2)+F_{2}(L-1m)=0"

"F_{2}(L-1m)=mg(L\/2)"

"F_{2}=mg(\frac{L}{2(L-1)})=20kg\times9.8m/s^{2}\times \frac{4m}{2(4m-1m)} = 130.66N"

"=131N"